# Transformers can do Bayesian Clustering

### v1

@@ 1 Introduction
Bayesian approaches address this limitation by placing prior distributions over model parameters, allowing for more robust and uncertainty-aware clustering. While this provides a principled way to incorporate uncertainty into the clustering process, inference in Bayesian models is typically computationally intensive, even when using approximate inference methods such as Variational Inference (VI) (Hoffman et al., 2013; Jordan et al., 1999; Wainwright & Jordan, 2008).
@@ 2 Background
To make inference tractable, several algorithms can be used, such as Variational Inference (VI) and Markov-Chain Monte Carlo (MCMC) (Andrieu et al., 2003; Neal, 2012; Welling & Teh, 2011). We focus on VI, which is commonly used for Bayesian clustering with GMMs (when using conjugate priors (Bishop, 2006, p. 474)). VI approximates the posterior with a simpler, factorized distribution: $q(\theta,\mathbf{z})=q(\theta)q(\mathbf{z})$ (Bishop, 2006; Blei & Jordan, 2006). This is known as the mean field approximation for the Bayesian Gaussian Mixture Model. As this assumption imposes restrictions on the form of these distributions, the variational distribution does not necessarily converge to the true posterior. Instead, it converges to the best possible approximation within the constraints of the variational family.
@@ 4 Experimental Setup
For hard cluster assignments, we employ three standard external metrics: Adjusted Rand Index (ARI) (Hubert & Arabie, 1985a), Adjusted Mutual Information (AMI) (Vinh et al., 2010), and Purity (Manning et al., 2008). ARI quantifies the agreement between predicted and true labels, adjusted for chance. AMI measures the mutual dependence between predicted and true labels, also adjusted for chance. Purity assesses the extent to which predicted clusters contain points from a single ground-truth class. To evaluate the probabilistic predictions, we compute the Negative Log-Likelihood (NLL) of the true cluster assignments under the model’s predicted posterior distribution. Since the cluster labels produced by the models may be permuted, we compute the NLL under all possible label permutations and report the minimum value obtained. Additional details of these metrics can be found in Appendix D.
@@
Among the Cluster-PFN models trained on the Easy and Hard priors, the one trained on the Hard prior performed best on real-world benchmarks. We evaluated it against baseline methods, applying feature-wise mean or median imputation on missing values (reporting only median results, as both gave similar outcomes) and using handcrafted model selection to assign cluster labels. VI was provided with the same priors used for the hard datasets.
@@ 5 Results
We begin by qualitatively analyzing Cluster-PFN’s clustering on the Old Faithful dataset, a classic real-world dataset. It records eruption durations and waiting times (Corduneanu & Bishop, 2001), and we use it to visually assess the Cluster-PFN’s performance. Figure 3 shows the clustering results produced by Cluster-PFN trained on the Easy prior on the Old Faithful dataset. The model assigns the highest posterior probability to 2 clusters, followed by 3 clusters. The partitionings for the clustering conditioned on 2 and 3 clusters are also shown, demonstrating that both align well with the underlying structure of the data.
@@ 8 Reproducibility statement
To ensure reproducibility, we fixed random seeds throughout all stages of our experiments. This includes synthetic data generation, evaluation of the models, as well as the creation of missingness masks, ensuring that results can be consistently replicated.
@@ C.1 Architecture
We use a cosine annealing (Loshchilov & Hutter, 2016) learning rate schedule with a warm-up, applied to an AdamW optimizer initialized with a tuned learning rate of 0.001.
@@ D.1 Adjusted Rand Index
Intuitively, the Rand Index evaluates the proportion of decisions (about whether a pair of points should be in the same cluster or not) that the clustering algorithm got correct. However, since the RI does not account for chance groupings, the Adjusted Rand Index introduces a normalization that adjusts the score for random labelings (Hubert & Arabie, 1985b).

### v2

@@ 1 Introduction
Bayesian approaches address this limitation by placing prior distributions over model parameters, allowing for more robust and uncertainty-aware clustering. While this provides a principled way to incorporate uncertainty into the clustering process, inference in Bayesian models is typically computationally intensive, even when using approximate inference methods such as Variational Inference (VI) (Hoffman et al., 2013; Jordan et al., 1999; Wainwright and Jordan, 2008).
@@ 2 Background
To make inference tractable, several algorithms can be used, such as Variational Inference (VI) and Markov-Chain Monte Carlo (MCMC) (Andrieu et al., 2003; Neal, 2012; Welling and Teh, 2011). We focus on VI, which is commonly used for Bayesian clustering with GMMs (when using conjugate priors (Bishop, 2006, p. 474)). VI approximates the posterior with a simpler, factorized distribution: $q(\theta,\mathbf{z})=q(\theta)q(\mathbf{z})$ (Bishop, 2006; Blei and Jordan, 2006). This is known as the mean field approximation for the Bayesian Gaussian Mixture Model. As this assumption imposes restrictions on the form of these distributions, the variational distribution does not necessarily converge to the true posterior. Instead, it converges to the best possible approximation within the constraints of the variational family.
@@ 4 Experimental Setup
For hard cluster assignments, we employ three standard external metrics: Adjusted Rand Index (ARI) (Hubert and Arabie, 1985a), Adjusted Mutual Information (AMI) (Vinh et al., 2010), and Purity (Manning et al., 2008). ARI quantifies the agreement between predicted and true labels, adjusted for chance. AMI measures the mutual dependence between predicted and true labels, also adjusted for chance. Purity assesses the extent to which predicted clusters contain points from a single ground-truth class. To evaluate the probabilistic predictions, we compute the Negative Log-Likelihood (NLL) of the true cluster assignments under the model’s predicted posterior distribution. Since the cluster labels produced by the models may be permuted, we compute the NLL under all possible label permutations and report the minimum value obtained. Additional details of these metrics can be found in Appendix D.
@@
Among the Cluster-PFN models trained on the Easy and Hard priors, the one trained on the Hard prior performed best on real-world benchmarks. We evaluated it against baseline methods, applying feature-wise mean or median imputation on missing values (reporting only median results, as both gave similar outcomes) and using handcrafted model selection to assign cluster labels. VI was provided with the same priors used for the hard datasets. S
@@ 5 Results
We begin by qualitatively analyzing Cluster-PFN’s clustering on the Old Faithful dataset, a classic real-world dataset. It records eruption durations and waiting times (Corduneanu and Bishop, 2001), and we use it to visually assess the Cluster-PFN’s performance. Figure 3 shows the clustering results produced by Cluster-PFN trained on the Easy prior on the Old Faithful dataset. The model assigns the highest posterior probability to 2 clusters, followed by 3 clusters. The partitionings for the clustering conditioned on 2 and 3 clusters are also shown, demonstrating that both align well with the underlying structure of the data.
@@ 8 Reproducibility statement
To ensure reproducibility, we fixed random seeds throughout all stages of our experiments. This includes synthetic data generation, evaluation of the models, as well as the creation of missingness masks, ensuring that results can be consistently replicated.Supplementary Material.
The supplementary material provides the implementation used in the experiments, along with the trained model weights.
@@ C.1 Architecture
We use a cosine annealing (Loshchilov and Hutter, 2016) learning rate schedule with a warm-up, applied to an AdamW optimizer initialized with a tuned learning rate of 0.001.
@@ D.1 Adjusted Rand Index
Intuitively, the Rand Index evaluates the proportion of decisions (about whether a pair of points should be in the same cluster or not) that the clustering algorithm got correct. However, since the RI does not account for chance groupings, the Adjusted Rand Index introduces a normalization that adjusts the score for random labelings (Hubert and Arabie, 1985b).